[Jeff, Daniel]
We added a 6" to 2.75" Zero Length Reducer Flange to a Laser Filter Cavity (LFC) 6" Gate Valve. I tightened the bolts to 34 Nm until the copper flange wasn't visible. I placed on the optics table a t-slot framing (80/20) structure with some viton to support the ion pump. We then placed the ion pump on the reducer flange and I tightened the bolts until the copper flange wasn't visible.
The cable to the ion pump fits, but it might constrain the composite 2.75" CF Angle Valve (that would go to the leak checker) from opening fully.
I would like to start on the 1550 path of OFC3. There are two ways to do this. We could just set up a test path directly from a fiber. With the intracavity alignment already done this seems like it is not worth the time. This just leaves aligning the light coming out of OFC2. The problem with this is as follows:
You can't PDH on 1550 in OFC2 as this was never set up. This means we have to lock on 775 and adjust the aom frequency so they are coresonant. This is bad for two reason: 1) the lock quality is very poor and 2) the two wavelengths are drifting relative to each other. You can fix this by adjusting the frequency of shift on the 775 path so that they are realigned, but continously doing so would be a pain. If you lock them while coresonant, and just watch the 1550 output over time it fluctuates wildly on a seconds time scale. I interpret this as the 1550 light drifting where it is along the lorentzian output peak. It does not seem feasible to try and align a cavity with this as your input light.
An alternate plan could be we fiber couple every cavity together. Use constant light from a real source to align and then just plug in your filter cavity output light fiber once its aligned. This introduced more loss into the GQuEST readout though. If anyone has ideas on this it would be appreciated.
Continuation of 12198.
I removed the lens in transmission of the cavity. This allowed me to see it was cleary clipping something circular. I walked the input beam and intracavity mirrors until the clipping was no longer observed. With this clipping gone was able to tune the alignment by eye on the camera a little more. At this point it is much better to use PDs to fine tune the alignment. I installed a 50:50 beam splitter in transmission as well as a PD. Additionally, I installed 2 mirrors in reflection of the cavity that act as steering mirrors onto a REFL PD. Next I aligned onto the REFL PD to allow for locking. While the cavity is locked, you can see the TRANS beam on the card to allow for easier alignment onto the TRANS PD.
775 path is fully installed and aligned. A cavity scan shows the cavity is well aligned. While locked, it seems there is some periodic signal on the error signal (this is laser locked with the teraxion) that is causing some drops in the amount of transmitted light power. By changing the coupling on REFL to DC you can check the visibility on the REFL PD. It seems pretty good, but may be improved by a lens as the diode is very small on these fast PDs used for PDH.
[Jeff, Daniel]
We unboxed the Agilent IDP 7 scroll pump and placed it on a platform meant to hold it. We placed this in the B111B Mobile Clean rooms close to its final location. We will need a 4 ft long KF25 hose (MH-QF-B48 on Kurt J Lesker is the cheapest option; we don't need additional features) to extend the included hose from the TwisTorr 74 turbo pump to the IDP 7. We have a 4 ft long 2.75" CF hose that's meant to be used in case we need to move the turbo off the LFC for vibration purposes (shouldn't be an issue if we use the ion pump to maintain vacuum), but the cost of two converters is about the same so we should get a single hose.
We also unboxed a lot of the other small components like fans, power cords, and the turbo controller. I forgot to install the inlet screen for the turbo pump, so I should take off the turbo, add it, and put the turbo back.
[Jeff, Daniel]
We put away the 75:25 fiber beam splitter, the FPC562 - Fiber Polarization Controller, some small items, and the breadboard that was on the B111B mobile clean room shelf. The 75:25 fiber beam splitter is in the bottom of the lista cabinet by the mobile clean rooms and the FPC562 is in the white drawers on the south wall (in a labeled drawer). We moved the now-empty breadboard into B102 by the other breadboards.
[Torrey, Jeff, Daniel]
We installed a second 6" CF Gate Valve onto the Laser Filter Cavity (LFC) facing the output cube. This is for the ion pump. Tightening the bolts was challenging with all of the other objects in the way. I don't think I maintained a consistent torque at the end of the tightening procedure as I couldn't fit the torque wrench in its torque limiting configuration. The bottom was showing more of the copper gasket than the top, so I started applying a bit more force. I believe I've tightened the bolts sufficiently and consistently enough to avoid leaks.
Since this is for the ion pump, I figured I should look more closely into what ion pumps, controllers, and cables we have from the Holometer. I found what I believe to be some controllers but I can't identify them. Based on a printed ion pump trouble shooting guide, the pumps appear to be from Duniway. The pumps we have appear to be the 30 L/s pumps (a bit tricky to identify from size but clearer from their weight and flange size), maybe the DSD-030-5125-M. I think the guide is from 2010. I also found a single cable that should go from the controller to the ion pump.
[Jeff, Torrey, Daniel]
We cleaned installed a 6" gate valve from the Holometer onto the top of the 6 way cross pointing towards the input. I placed the handle to be more accessible from the side of the table. I tightened the bolts by hand until there was steel to steel contact. I couldn't fit the torque wrench.
I then added a 6" to 4.5" reducer flange on top of the gate valve and tightened the bolts to 34 Nm until there was steel to steel contact all around.
We then added the Agilent TwisTorr 74 Turbo Pump onto the reducer flange. We could not fit screws when they had washers attached, so we did not use any washers. There is *very* little tolerance for getting a wrench around the screw heads such that we had to center the pump more precisely than usual. Even with that, we could fit only the torque adapter wrench around the screw (and barely at that). I think we should buy another (not another set) that I'll grind/cut down to be open face/nearly open face for next time. I didn't find any 5/16" ID slotted washers on McMaster with small enough OD. I tightened the screws to 20 Nm as per the manual. The copper was still visible. It should be easy to determine if there is a leak here (although it would be indistinguishable from one within the pumping system) by shutting the gate valve.
I've started alignment of OFC3. I've decided to do 775 path first. It is aligned to the point where I see first pass light (i.e. passes through M1 and M2 and onto camera). As well as the second pass light interfering with the first pass to create fringes. In the past once I've been at this point, you scan around on the intercavity mirrors until you see the third pass beam, and then turn on some kind of cavity length/frequency scan to see flashes. I haven't found the third pass beam yet. Hopefully should have a third, functioning 775 cavity soon.
One note: In order for the reflected beam incident on the cavity to clear the input mirror I had to move the two input mirrors back about 1.5 inches each. I also rotated the mount for the second input mirror to give additional clearance. The REFL beam clears the input mirror but only by about an inch or less.
I added a f=100mm lens to the TRANS output and moved the camera back. Then realigned without the fourth mirror on, put the fourth mirror back on and a scan on one of the mirrors revealed the third pass beam. It is now touched up enough to see some 1,0 0,1 and 0,0s flashing. At this point I find a PD is better for alignment. Will set that up after journal club.
Logging an error message that popped up when using the LXM Control software. Unsure what it means. Mainly a reminder to myself to check the manual later.
I cleaned some 1/4-20 button head screws, 1/2" long with acetone and isopropanol since they were greasy. I used these and some spring tab t-nuts that don't slide down under gravity to reassemble the 80/20 structure to support the bottom and sides of the 8" CF to 6" CF reducing tee on the Laser Filter Cavity (LFC) Vacuum System. To be honest, I'm not sure how much it's actually helping support the weight. I'm also using a scissor stand platform to help with the weight. I did not assembled the 45° brackets since it will be easier to install them with more spring tab t-nuts.
Note for others when assembling 80/20: a small hex key is great at positioning the t-nuts, especially the cheap ones that can slide around.
Note for me and others going forward: thicker viton (~1/4" thick) would make the contact between the 80/20 and the tee much better.
[Sander, Daniel]
I removed a 2.75" CF Flange Tee from a Holometer ion pump sub-system and cleaned tape residue from it. I then installed it on the North 6" to 2.75" zero length reducer flange (it is easily accessible from the side of the table). Sander and I cleaned and installed two 2.75" angle valves onto the tee, one on each side. The all metal angle valve is for the SRS RGA 100 (which I unboxed and inspected; it looks good) to reduce any contamination. The composite angle valve is for a leak checker. I tightened all 3 joints by hand unitl no copper was visible and there was a steel to steel connection all the way around.
I tested both valves today. The all metal angle valve requires a 1/2" square drive. It took ~30 Nm of torque to loosen it. I turned the valve a few turns and then turned it back. I also loosened and tightened the composite angle valve.
The all metal angle valve is from the power recycling cube. I covered that hole in UHV foil.
I have now set up the particle counter and a NUC to run the software I made to view and display the current data coming from the counter. We will however need to continue to turn the counter back on to keep counting if it stops running. We should be able to tell via the interface i made if it stopped (past 6 hours or other plots are blank).
Anyone on the subnet can access this page if it is running, see the pdf's below for how it should look.
Access particle counter page: http://192.168.56.1:8050
Access the particle counter data page: http://192.168.248.101/
(Both are only available on the subnet)
Notes:
I installed the Up-to-air Valve (F0275XVALVE) on the "input" side of the Laser Filter Cavity (LFC) on a 6" to 2.75" reducer flange. I tightened the bolts to ~20 Nm but needed to use a basic wrench while tightening the bolts.
[Jeff, Alex, Daniel]
We placed the partially assembled 6" 6 Way Cross on the 6" flange of the 8" to 6" reducer tee of the Laser Filter Cavity (LFC). The 2" long screws have to go into the 6 way cross, so we put them through the rotatable flange on the reducer tee. The four 2" long screws that go over the body of the tee must be inserted into the tee's flange alone, rotated into place, then inserted into the 6 way cross. We then added the remaining screws and plate nuts. The 80/20 structure got in the way so I partially disassembled it and instead and supporting the tee with a small jack and some viton. I tightened the screws to 34 Nm until the copper gasket wasn't visible.
We then added two 6" to 2.75" zero length reducer flanges, one on the "input" side of the LFC (a non-rotatable flange for the 6 way cross) and one on the open end of the table away from the power distribution. I need to tighten these screws and will then add the subcomponents that go on the 2.75" side. The subcomponent on the open end of the table is a 2.75" tee that I sourced from the holometer. I need to separate it from the ion pump and valve.
Jeff and I made sure to orient the 6" to 2.75" zero length reducer flanges so that everything looks vertical/horizontal instead of askew.
I tightened the flanges to 34 Nm and the copper gaskets weren't visible.
[Jeff, Daniel]
We assembled a 6" to 2.75" zero length reducer flange on a non-rotatable flange of the 6" 6 way cross for the Laser Filter Cavity (LFC). The 6 way cross has 3 rotatable flanges and 3 non-rotatable flanges. Opposite sides have one of each.
After tightening the bolts to 34 Nm and seeing no copper gasket, we added the Agilent FRG702 Pressure Gauge. The magnet was 0.6" below the main body of the gauge before moving it up to fit the screws. We added the screws and tightened as hard as I could with the ~8" long wrench. The copper gasket wasn't visible. We tried to move the magnet down to its original location but could only get it 0.45" in down and unsecured.
We tried to add the 6 way cross to the 8" to 6" reducer tee but it was too heavy for 2 people if one person help up the rotatable flange on the tee. We'll need a 3rd person for this job.
I aligned the magnet's rotation as shown in the attached picture from my Agilent contact. I also moved the magnet away from the body of the pressure gauge, but apparently this degree of freedom doesn't matter as much. I did find a pressure gauge that works down to 10^-4 Torr from the Holometer. Maybe we can attach that and use it as a check.
Now that we have an operational timing circuit, we can assemble the main chassis of the Babbage cymac and hopefully not neet to poke around inside anymore. The lid does not fit on top due to the bulk of the pcie extension cables, so Daniel is going to cut a hole in the top and the cables will poke out.
Notes from assembly:
- The standoff holes on the General Standards ADC and DAC are a bit small for the 6-32 standoffs. However the standoffs are nylon and can be forced through.
- The angle brackets for holding the Adnacom backplane in place are not the correct size. Daniel made something work but it required stacking 7 washers to make a spacer.
- The main headache of this assembly are the pcie extension cables. Most of the cables Todd and I can find on Amazon are quite bulky, and there is very little room to fit them between the PCBs and the walls of the chassis. I ordered a few types and tried a few of Todd's, and they are all bulky. Todd has some old ones that work well but he can't find them sold online anymore. Instead of continuing to deal with this issue we are cutting a hole in the chassis.
Big todos on this project are to order the input/output PCBs (for whitening, dewhitening, and possibly other tasks?), and to push the software along so that we can begin communicating with the ADCs and DACs.
The assembled chassis with part of the lid removed to accomodate the pcie cables.